light travels (in vacuum) at a velocity of 3 x 10 ^8 m/s. The lowest visible wavelength is 420 nanometers ( 1 nanometer = 1 x 10^-9) Calculate the corresponding frequency for that wavelength.
Answer #1

We know, frequency of light wave = speed of light in vaccum / wavelength

That is ν = c / λ

Therefore for wavelength,λ = 420nm = 420 * 10^-9m and speed of light,c = 3* 10^8m/s,

We have ν= (3*10^8 m/s) / (420*10^-9 m)

ν = 7.143 * 10^14 Hertz. Answer
Answer #2

 Calculation:

1. Key Formula:

The speed of light (c), wavelength (λ), and frequency (f) are related by:

c=λ×forf=cλ

  • Speed of light (c): 3×108m/s

  • Wavelength (λ): 420nm (convert to meters first!).


2. Convert Wavelength to Meters:

1nm=109m    420nm=420×109m=4.2×107m

3. Plug into the Formula:

f=cλ=3×108m/s4.2×107m=34.2×108+7Hzf=0.714×1015Hz=7.14×1014Hz

4. Simplify (Optional):

  • 7.14×1014Hz=714THz (terahertz).

 
  1. Answer: 7.14 × 10¹⁴ Hz.


answered by: anonymous